# A quasilinear Keller-Segel model with saturated discontinuous advection

### v1

@@ 1. Introduction
and perform nonlinear stability analysis. In particular they show that if the initial data is near, in some norm, to the equilibrium function $\rho_{\infty}=\frac{\chi}{2}e^{-\chi|x-x_{0}|}$, then the corresponding solution converges to $\rho_{\infty}$ exponentially fast.
@@
$\left\{\begin{array}[]{rcll}\partial_{t}\rho-\textrm{div}\,\left[\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\right]&=&0,&\quad\mbox{in }\mathbb{R}^{d}\times(0,\infty),\\ \vspace{2mm}\rho(0)&=&\rho_{in},&\quad\mbox{in }\mathbb{R}^{d},\end{array}\right.$ (9)
@@
This monotonicity property, coupled with the theory of monotone operators and a fixed-point argument, proves fundamental in the proof of existence of solutions to (9) equipped with an extra viscosity term $\varepsilon\Delta\rho$. The crux of the analysis is the vanishing viscosity limit $\varepsilon\to 0$ and the corresponding identification of the limit of the nonlinear term
@@
[…]
and
Here is our main result:
@@ 2. A priori estimates and technical results
for any $\varphi\in L^{2}(0,T,H^{1}(\mathbb{R}^{d}))$ with initial smooth data $\rho_{0}$. The function $\rho_{0}$ belongs to $W^{1,p}\cap L^{\infty}\cap L^{1}_{2}(\mathbb{R}^{d})$, with $p>d$.
The following a priori estimates hold:
@@
By monotone convergence we can take the limit $R\rightarrow\infty$ and conclude that $\int\rho_{\varepsilon}\leq\int\rho_{0}$. Now that we know that the mass of $\rho_{\varepsilon}$ is bounded, we test again (11) with $\varphi$ , but this time we pass directly to the limit $R\rightarrow\infty$ in (12) to find
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\rho_{\varepsilon}\,{\rm d}x)(T)+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\lvert\nabla\rho_{\varepsilon}\rvert^{2}}{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t\leq\bar{C}[\rho_{0}]$
@@
where $C[\rho_{0}]$ is a constant that depends on $\int_{\mathbb{R}^{d}}\rho_{0}|\log{\rho_{0}}|\,{\rm d}x$, the mass and the second moment of $\rho_{0}$. To bound $\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}$, we define $h[\rho_{\varepsilon}](t):=\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log(\rho_{\varepsilon})\,{\rm d}x\right)(t)$. Using the fact that
@@
with $c$ independent on $\varepsilon$. Summarizing, $\textrm{div}(u)\in L^{p}(0,T,W^{1,p})$, $g\in L^{\infty}(t,T,W^{1,p})$, and the thesis follows. ∎
@@
where $\beta=1+\frac{2}{d+2}$, using (16) to obtain the second and the third inequality. From [9, Lemma 7.1] it follows that $U_{k}\rightarrow 0$ as $k\rightarrow\infty$, i.e. that $\rho_{\varepsilon}\leq M$, if
@@
Let $\{f_{n}\}$ be a bounded sequence in $L^{2}(0,T;H^{1}(\mathbb{R}^{d}))\cap L^{2}(0,T;\mathbb{R}^{d},\sqrt{1+|x|^{2}}\,{\rm d}x)$ such that $\partial_{t}f_{n}$ is bounded in $L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))$, then
@@ 3.1. Existence for the semi-linearized equation
$\forall\varphi\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$, where $\rho_{\varepsilon}(0,\cdot)=\rho_{{\rm in},\varepsilon}$ and $\rho_{{\rm in},\varepsilon}$ is a $W^{1,p}\cap L^{\infty}(\mathbb{R}^{d})$ approximation of $\rho_{{\rm in}}$ such that $\rho_{{\rm in},\varepsilon}\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ when $\varepsilon\rightarrow 0$. In order to prove the existence of the above equation, we use general theory of monotone operators, concretely, [18, Theorem 30.A]. We define the following operator $A$ as
@@
which must satisfy the following properties: 1. Monotonicity. . The operator $A$ is monotone, since $\forall\;u,v\in H^{1}(\mathbb{R}^{d})$ we have
@@
by dominated convergence. 3. Coerciveness. We just compute
@@
so $A$ is coercive. 4. Growth condition. For this we have to prove that $\|A[u]\|_{H^{*}(\mathbb{R}^{d})}\leq C\|u\|_{H(\mathbb{R}^{d})}$, but, by Hölder,
@@
5. Measurability. By assumption. Then, by [18, Theorem 30.A], there exist a unique weak solution $\rho_{\varepsilon}\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$ to (3.1) with $\partial_{t}\rho\in L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))$ satisfying the initial condition in the sense of the strong $L^{2}(\mathbb{R}^{d})$ convergence:
@@ 3.2. Existence for approximated solutions
Now, we are going to use the Leray-Schauder fixed point theorem to show the existence of a solution to (11). To use the theorem we require $\mathcal{T}$ to be compact. First, using $\rho_{\varepsilon}$ as test function in (3.1) we get
@@
By Lemma 4, $\mathcal{T}$ maps bounded sets to precompact sets. Now, we have to prove that the map is continuous, i.e. that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma_{n},z_{n})=\mathcal{T}(\sigma,z)$ when $z_{n}\rightarrow z$ strongly in $L^{2}(0,T,L^{2}(\mathbb{R}^{d}))$ and $\sigma_{n}\rightarrow\sigma$ in $[0,1]$. Is easy to see that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma_{n},z_{n})=\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma,z_{n})$. Then, as the map is compact, it only remains to prove
@@
We conclude that $\{\nabla\rho_{\varepsilon,n}\}$ is a Cauchy sequence in $L^{2}(0,T,L^{2}(\mathbb{R}^{d})))$, and therefore, $\nabla\rho_{\varepsilon,n}\rightarrow\nabla\rho_{\varepsilon}$ a.e. $x\in[0,T]\times\mathbb{R}^{d}$. The map $\mathcal{T}$ is continuous, compact and constant when $\sigma=0$. Moreover, fixed points are uniformly bounded in $\sigma$. Finally, we have to prove that fixed points of $\mathcal{T}$ are nonnegatives so we can remove the $+$ sign. Using Stampacchia’s truncation argument, we introduce $(\rho_{\varepsilon})_{-}$ in (11) as test function and we get
@@
If $\rho_{{\rm in},\varepsilon}\geq 0$, solutions to $\mathcal{T}(\sigma,\rho_{\varepsilon})=\rho_{\varepsilon}$ are nonnegatives. Then, Leray-Schauder fixed-point yields a solution to
@@ 4. The limit $\varepsilon\rightarrow 0$ and proof of Theorem 1
The identification $G=\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}\nabla\rho$ will follow by almost everywhere convergence of $\rho_{\varepsilon}$ (given by Lemma (4)) and of $\nabla\rho_{\varepsilon}$. The next two subsections will be devoted to proof that
@@ 4.1. Entropy method
The relative entropy production $\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]$ is nonnegative as a direct consequence of the monotonicity of the operator $\left(1-\frac{\chi}{|\nabla\log\rho|}\right)_{+}\nabla\log\rho$. The error terms originate from the viscosities $\varepsilon\Delta\rho_{\varepsilon}$ and $\delta\Delta\rho_{\delta}$ As $\varepsilon,\delta\rightarrow 0$, $H[\rho_{{\rm in},\varepsilon}|\rho_{{\rm in},\delta}]\rightarrow 0$ by construction, and the error terms also vanish. Since both $H$ and $\mathcal{D}$ are nonnegative, as $\delta,\varepsilon\to 0$, we must have
@@
where $E_{\varepsilon,\delta}$ are error terms, which must vanish when $\sigma\rightarrow 0$.
Note that (26) is an approximation of $\int_{0}^{T}\frac{d}{\,{\rm d}t}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}]$. We will show
@@
${I_{3}}:=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t$ (35)
@@ 4.4. Obtaining the initial data $\rho_{{\rm in}}$ for $\rho$
In what follows, we will show that $\rho(t)\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$.
We start by reminding that $\rho_{\varepsilon}$ is bounded in the space
@@
which embeds continuously into $C^{0}([0,T];L^{2}(\mathbb{R}^{d}))$ [17, Prop. 23.23]. In particular it follows that $\rho_{\varepsilon}(t)\rightharpoonup\rho(t)$ weakly in $H^{-1}(\mathbb{R}^{d})$ for every $t\in[0,T]$, hence
@@
$\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}\leq\liminf_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\limsup_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}$
@@ 4.5. Uniqueness
$=\int_{\mathbb{R}^{d}}(\rho_{2}+\sigma)\left[\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)+1\right](T)\,{\rm d}x$

### v2

@@ Abstract.
Key words and phrases:
Degenerate diffusion, parabolic PDEs, Keller-Segel, existence of weak solutions, quasilinear PDEs
@@ 1. Introduction and main result
and perform nonlinear stability analysis for the one dimensional case. In particular they show that if the initial data is near, in some norm, to the equilibrium function $\rho_{\infty}=\frac{\chi}{2}e^{-\chi|x-x_{0}|}$, then the corresponding solution converges to $\rho_{\infty}$ exponentially fast. The constant $\alpha\geq 0$ denotes the natural decay of the chemical component.
@@
$\left\{\begin{array}[]{rcll}\partial_{t}\rho-\textrm{div}\,\left[\left(1-\frac{\chi\rho}{\lvert\nabla\rho\rvert}\right)_{+}\nabla\rho\right]&=&0,&\quad\mbox{in }\mathbb{R}^{d}\times(0,\infty),\vspace{2mm}\\ \vspace{2mm}\rho(0)&=&\rho_{in},&\quad\mbox{in }\mathbb{R}^{d},\end{array}\right.$ (9)
@@
This monotonicity property, coupled with the theory of monotone operators and a fixed-point argument, proves fundamental in the proof of existence of solutions to (9) equipped with an extra viscosity term $\varepsilon\Delta\rho$. The crux of the analysis will be the vanishing viscosity limit $\varepsilon\to 0$ and the corresponding identification of the limit of the nonlinear term
@@
Before we dwell with analytical details, let us mention that a kinetic formulation and related stability analysis of (3) in one dimensional space has been considered in several works [18, 8, 9, 3]. The singular limit $\alpha\to+\infty$ in this setting is still an open problem.
Our main results are summarized in the following theorem:
@@
As mentioned earlier, the most important observation is that the diffusion operator in (9) is degenerate, but monotone. The construction of solution to (9) with extra viscosity and a-priori estimates uniform in the viscosity parameter will follows automatically from the theory of monotone operators. The crux of the analysis is, however, the identification of the limit of the nonlinear term in the vanishing viscosity limit:
[…]
and
@@ 2. A priori estimates and technical results
for any $\varphi\in L^{2}(0,T,H^{1}(\mathbb{R}^{d}))$ with initial smooth data $\rho_{0}$. The function $\rho_{0}$ belongs to $W^{1,p}\cap L^{\infty}\cap L^{1}_{2}(\mathbb{R}^{d})$, with $p>d$. The following a priori estimates hold:
@@
By monotone convergence we can take the limit $R\rightarrow\infty$ and conclude that $\int\rho_{\varepsilon}\leq\int\rho_{0}$.
Now that we know that the mass of $\rho_{\varepsilon}$ is bounded, we test again (11) with $\varphi$ , but this time we pass directly to the limit $R\rightarrow\infty$ in (12) to find
@@
$(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log\rho_{\varepsilon}\,{\rm d}x)(T)+\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{\lvert\nabla\rho_{\varepsilon}\rvert^{2}}{\rho_{\varepsilon}}\,{\rm d}x\,{\rm d}t\leq\bar{C}[\rho_{0}],$
@@
where $C[\rho_{0}]$ is a constant that depends on $\int_{\mathbb{R}^{d}}\rho_{0}|\log{\rho_{0}}|\,{\rm d}x$, the mass and the second moment of $\rho_{0}$.
To bound $\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{|\nabla\rho_{\varepsilon}|^{2}}{\rho_{\varepsilon}}$, we define $h[\rho_{\varepsilon}](t):=\left(\int_{\mathbb{R}^{d}}\rho_{\varepsilon}\log(\rho_{\varepsilon})\,{\rm d}x\right)(t)$. Using the fact that
@@
with $c$ independent on $\varepsilon$. Summarizing, $\textrm{div}(u)\in L^{p}(0,T,W^{1,p})$, $g\in L^{\infty}(t,T,W^{1,p})$, and the thesis follows.
@@
where $\beta=1+\frac{2}{d+2}$, using (16) to obtain the second and the third inequality. From [12, Lemma 7.1] it follows that $U_{k}\rightarrow 0$ as $k\rightarrow\infty$, i.e. that $\rho_{\varepsilon}\leq M$, if
@@
Let $\{f_{n}\}$ be a bounded sequence that belongs to $L^{2}(0,T;H^{1}(\mathbb{R}^{d}))\cap L^{2}(0,T;\mathbb{R}^{d},\sqrt{1+|x|^{2}}\,{\rm d}x)$ and such that $\partial_{t}f_{n}$ is bounded in $L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))$. We have
@@ 3.1. Existence for the semi-linearized equation
$\forall\varphi\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$, where $\rho_{\varepsilon}(0,\cdot)=\rho_{{\rm in},\varepsilon}$ and $\rho_{{\rm in},\varepsilon}$ is a $W^{1,p}\cap L^{\infty}(\mathbb{R}^{d})$ approximation of $\rho_{{\rm in}}$ such that $\rho_{{\rm in},\varepsilon}\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ when $\varepsilon\rightarrow 0$. In order to prove the existence of the above equation, we use general theory of monotone operators, concretely, [22, Theorem 30.A]. We define the following operator $A$ as
@@
which must satisfy the following properties:
1. Monotonicity. . The operator $A$ is monotone, since $\forall\;u,v\in H^{1}(\mathbb{R}^{d})$ we have
@@
by dominated convergence.
3. Coerciveness. We just compute
@@
so $A$ is coercive.
4. Growth condition. For this we have to prove that $\|A[u]\|_{H^{*}(\mathbb{R}^{d})}\leq C\|u\|_{H(\mathbb{R}^{d})}$, but, by Hölder,
@@
5. Measurability. By assumption.
Then, by [22, Theorem 30.A], there exists a unique weak solution $\rho_{\varepsilon}\in L^{2}(0,T;H^{1}(\mathbb{R}^{d}))$ to (3.1) with $\partial_{t}\rho\in L^{2}(0,T;H^{-1}(\mathbb{R}^{d}))$ satisfying the initial condition in the sense of the strong $L^{2}(\mathbb{R}^{d})$ convergence:
@@ 3.2. Existence for approximated solutions
Now, we are going to use the Leray-Schauder fixed point theorem to show the existence of a solution to (11). Leray-Schauder theorem requires $\mathcal{T}$ to be compact. First, using $\rho_{\varepsilon}$ as test function in (3.1) we get
@@
By Lemma 4, $\mathcal{T}$ maps bounded sets to precompact sets.
Now, we have to prove that the map is continuous, i.e. that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma_{n},z_{n})=\mathcal{T}(\sigma,z)$ when $z_{n}\rightarrow z$ strongly in $L^{2}(0,T,L^{2}(\mathbb{R}^{d}))$ and $\sigma_{n}\rightarrow\sigma$ in $[0,1]$. Is easy to see that $\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma_{n},z_{n})=\lim\limits_{n\rightarrow\infty}\mathcal{T}(\sigma,z_{n})$. Then, as the map is compact, it only remains to prove
@@
We conclude that $\{\nabla\rho_{\varepsilon,n}\}$ is a Cauchy sequence in $L^{2}(0,T,L^{2}(\mathbb{R}^{d})))$, and therefore, $\nabla\rho_{\varepsilon,n}\rightarrow\nabla\rho_{\varepsilon}$ a.e. $x\in[0,T]\times\mathbb{R}^{d}$. The map $\mathcal{T}$ is continuous, compact and constant when $\sigma=0$. Moreover, fixed points are uniformly bounded in $\sigma$.
Finally, we have to prove that fixed points of $\mathcal{T}$ are nonnegatives so we can remove the $+$ sign. Using Stampacchia’s truncation argument, we introduce $(\rho_{\varepsilon})_{-}$ in (11) as test function and we get
@@
If $\rho_{{\rm in},\varepsilon}\geq 0$, solutions to $\mathcal{T}(\sigma,\rho_{\varepsilon})=\rho_{\varepsilon}$ are nonnegatives.
Then, Leray-Schauder fixed-point yields a solution to
@@ 4. The limit $\varepsilon\rightarrow 0$ and proof of Theorem 1
The identification $G=\left(1-\frac{\chi\rho}{|\nabla\rho|}\right)_{+}\nabla\rho$ will follow by almost everywhere convergence of $\rho_{\varepsilon}$ (given by Lemma (4)) and of $\nabla\rho_{\varepsilon}$. The next two subsections will be devoted to proving that
@@ 4.1. Entropy method
The relative entropy production $\mathcal{D}[\rho_{\varepsilon}|\rho_{\delta}]$ is nonnegative as a direct consequence of the monotonicity of the operator $\left(1-\frac{\chi}{|\nabla\log\rho|}\right)_{+}\nabla\log\rho$. The error terms originate from the viscosities $\varepsilon\Delta\rho_{\varepsilon}$ and $\delta\Delta\rho_{\delta}$. As $\varepsilon,\delta\rightarrow 0$, $H[\rho_{{\rm in},\varepsilon}|\rho_{{\rm in},\delta}]\rightarrow 0$ by construction, and the error terms also vanish. Since both $H$ and $\mathcal{D}$ are nonnegative, as $\delta,\varepsilon\to 0$, we must have
@@
where $E_{\varepsilon,\delta}$ are error terms, which must vanish when $\sigma\rightarrow 0$. Notice that (26) is an approximation of $\int_{0}^{T}\frac{d}{\,{\rm d}t}H_{{\rm rel}}[\rho_{\varepsilon}|\rho_{\delta}]$. We will show
@@
${I_{3}}:=-\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(1-\frac{\chi\rho_{\varepsilon}}{\lvert\nabla\rho_{\varepsilon}\rvert}\right)_{+}\nabla\rho_{\varepsilon}\nabla\left(\log\frac{\rho_{\varepsilon}+\sigma}{\rho_{\delta}+\sigma}\right)\phi_{R}\,{\rm d}x\,{\rm d}t,$ (35)
@@ 4.4. Obtaining the initial data $\rho_{{\rm in}}$ for $\rho$
In what follows, we will show that $\rho(t)\rightarrow\rho_{{\rm in}}$ strongly in $L^{p}(\mathbb{R}^{d})$ as $t\to 0$. We start by recalling that $\rho_{\varepsilon}$ is bounded in the space
@@
which embeds continuously into $C^{0}([0,T];L^{2}(\mathbb{R}^{d}))$ [21, Prop. 23.23]. In particular it follows that $\rho_{\varepsilon}(t)\rightharpoonup\rho(t)$ weakly in $H^{-1}(\mathbb{R}^{d})$ for every $t\in[0,T]$, hence
@@
$\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})}\leq\liminf_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\limsup_{t\to 0}\|\rho(t)\|_{L^{p}(\mathbb{R}^{d})}\leq\|\rho_{0}\|_{L^{p}(\mathbb{R}^{d})},$
@@ 4.5. Uniqueness
$=\int_{\mathbb{R}^{d}}(\rho_{2}+\sigma)\left[\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)\log{\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)}-\left(\frac{\rho_{1}+\sigma}{\rho_{2}+\sigma}\right)+1\right](T)\,{\rm d}x.$
@@
5. Disclosure statement
There are no relevant financial or non-financial competing interests to report.